# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v1

@@ Appendix I Constrained Convex Optimisation
$N_{h}^{t}(s,a)>\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\left(\frac{SAH}{\delta^{\prime}}\right),$
then
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{N_{h}^{t}(s_{h}^{t},a_{h}^{t})\vee 1}}\mid\mathcal{F}_{t-1}\right]\leq\tilde{\mathcal{O}}(\sqrt{SAH^{2}T}+SAH).$
@@
    $\big{|}\hat{l}_{h}^{t}(s,a)-l_{h}(s,a)\big{|}\lesssim\sqrt{\frac{1}{N_{h}^{t}(s,a)}}$.
@@
    $\big{|}\overline{p}_{h}^{t}(s^{\prime}\mid s,a)-p_{h}(s^{\prime}\mid s,a)\big{|}\lesssim\sqrt{\frac{p_{h}(s^{\prime}\mid s,a)}{N_{h}^{t}(s,a)\lor 1}}+\frac{1}{N_{h}^{t}(s,a)\lor 1}$.
@@
    $N_{h}^{t}(s,a)\leq\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{t}}(s,a\mid p)-H\ln\frac{SAH}{\delta^{\prime}}$.

### v2

@@ Appendix I Constrained Convex Optimisation
$N_{h}^{t-1}(s,a)>\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\left(\frac{SAH}{\delta^{\prime}}\right),$
then for all $T^{\prime}\leq T$,
$\sum_{t^{\prime}=1}^{T^{\prime}}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{N_{h}^{t^{\prime}-1}(s_{h}^{t^{\prime}},a_{h}^{t^{\prime}})\vee 1}}\mid\mathcal{F}_{t^{\prime}-1}\right]\leq\tilde{\mathcal{O}}(\sqrt{SAH^{2}T^{\prime}}+SAH).$
@@
    $\big{|}\overline{l}_{h}^{t}(s,a)-l_{h}(s,a)\big{|}\lesssim\sqrt{\frac{1}{N_{h}^{t-1}(s,a)}}$.
@@
    $\big{|}\overline{p}_{h}^{t}(s^{\prime}\mid s,a)-p_{h}(s^{\prime}\mid s,a)\big{|}\lesssim\sqrt{\frac{p_{h}(s^{\prime}\mid s,a)}{\max\{N_{h}^{t-1}(s,a),1\}}}+\frac{1}{\max\{N_{h}^{t-1}(s,a),1\}}$.
@@
    $N_{h}^{t-1}(s,a)\leq\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\frac{SAH}{\delta^{\prime}}$.